OpenStudy (anonymous):

If the diagonal if a square is (50-12 square root of 2) cm, what is the area of the square?

OpenStudy (anonymous):

@jiteshmeghwal9 pls help again :)) sorry to disturb you

OpenStudy (anonymous):

Diagonal of the square = Side x sqrt(2) Area of the square = (side)^2 These are the basic theorems we should use. Associating are of the square in terms of it's side, which is associated to the diagonal.

OpenStudy (anonymous):

D = 50 - 12sqrt(2) D = (s)sqrt(2) A = s^2 A = (50-12sqrt(2))^2/2

OpenStudy (anonymous):

\[A = {({50-12\sqrt{2})}^2}/2\]

OpenStudy (anonymous):

that cant be true if the side is\[50-12\sqrt{2}\]

OpenStudy (anonymous):

I mean diagonal is the diagonal \[5-12\sqrt{2}\] or \[(5-12)\sqrt{2}\]

OpenStudy (anonymous):

A = 545.47 cm^2

OpenStudy (anonymous):

If we consider \[50 - 12\sqrt{2}\] as the diagonal

OpenStudy (anonymous):

if it is\[50-12\sqrt{2}=(\frac{50}{\sqrt{2}}-12)\sqrt{2}\] side\[50/\sqrt{2}-12\] \[A=(\frac{ 50}{\sqrt{2}}-12)^2\]

OpenStudy (anonymous):

Still the same answer

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

A = 545.47 cm^2